OpenStudy (anonymous):

Solve for x: |x+15|=|x-2|

OpenStudy (anonymous):

there is a absolute value

OpenStudy (anonymous):

SOLVE TWO EQUATIONS x+15=x-2 and x+15=-(x-2)

OpenStudy (anonymous):